# Movable Antennas Meet Intelligent Reflecting Surface: Friends or Foes?

### v1

@@ V-B Multi-User System
Figure 6: Sum-rate under versus the BS-IRS distance under the LoS BS-IRS channel.
@@
Figure 8: Sum-rate under versus the BS-IRS distance under the multi-path BS-IRS channel.
@@ VI Conclusions and Future Directions
In this paper, we investigated a joint active/passive beamforming and BS antenna position optimization problem for an IRS-assisted multi-user MISO MA system. To gain useful insights, we first conducted performance analysis and solved the SNR maximization problem in the single-user case. Then, we solved the sum-rate maximization problem using the AO algorithm in the multi-user case. Both of our analytical and numerical results revealed that the presence of an IRS can hinder the performance gain of the MAs over conventional FPAs if the BS-user direct link is blocked. The main takeaways are summarized as follows.
  1. 1.
    For IRS-assisted MA systems, if the BS-IRS channel is LoS-dominated and the IRS is configured with an optimized reflection, the performance gain of MAs over FPAs diminishes with the BS-IRS distance and ultimately vanishes.
  2. 2.
    If the BS-IRS channel has non-negligible multi-path components, MAs generally yield a more significant gain over FPAs compared to the LoS-dominated BS-IRS channel, under both near- and far-field propagation conditions.
  3. 3.
    MAs tend to yield a more significant gain over FPAs if the IRS is configured with a random reflection instead of an optimized one.
This paper can be extended to various directions as future work, as listed below.
  1. 1.
    First, we only consider the cascaded BS-IRS-user channel, ignoring the direct BS-user channel. It would be interesting to further investigate how the performance gain varies with the strength ratio of the direct link to the reflected link.
  2. 2.
    Second, we only consider the BS-side MAs in this paper, and their positions need to cater to all users via only their shared BS-IRS channel. Hence, it is worthy of investigating the performance gain of user-side MAs over FPAs, which may have a larger DoF for performance enhancement by reconfiguring their IRS-user channels individually.
  3. 3.
    Last but not least, we assume perfect CSI on all links to focus on the performance analysis and optimization in this paper. However, the coexistence of the IRS and MAs may significantly increase the CSI estimation overhead. It is thus interesting to study whether these two technologies can be reciprocal in terms of channel estimation.

### v2

@@ V-B Multi-User System
Figure 6: Sum-rate versus the BS-IRS distance under the LoS BS-IRS channel.
@@
Figure 8: Sum-rate versus the BS-IRS distance under the multi-path BS-IRS channel.
@@ VI Conclusions and Future Directions
In this paper, we investigated a joint active/passive beamforming and BS antenna position optimization problem for an IRS-assisted multi-user MISO MA system. To gain useful insights, we first conducted performance analysis and solved the SNR maximization problem in the single-user case. Then, we solved the sum-rate maximization problem using the AO algorithm in the multi-user case. Both of our analytical and numerical results revealed that the presence of an IRS can hinder the performance gain of the MAs over conventional FPAs if the BS-user direct link is blocked. The main takeaways are summarized as follows. First, for IRS-assisted MA systems, if the BS-IRS channel is LoS-dominated and the IRS is configured with an optimized reflection, the performance gain of MAs over FPAs diminishes with the BS-IRS distance and ultimately vanishes. Second, if the BS-IRS channel has non-negligible multi-path components, MAs generally yield a more significant gain over FPAs compared to the LoS-dominated BS-IRS channel, under both near- and far-field propagation conditions. Third, the MAs tend to yield a more significant gain over FPAs if the IRS is configured with a random reflection instead of an optimized one.
This paper can be extended to various directions as future work. First, we only consider the cascaded BS-IRS-user channel, ignoring the direct BS-user channel. In the presence of the direct BS-user channel, the performance gain of MAs depends on more factors such as the strength ratio of the direct link to the reflected link, as well as the characteristics of the individual BS-user and BS-IRS channels, which renders the performance analysis more challenging and deserves further in-depth investigation. Second, we only consider BS-side MAs in this paper, and their positions need to cater to all users via only their shared BS-IRS channel. Hence, it is worthy of investigating the performance gain of user-side MAs over FPAs, which may have a larger DoF for performance enhancement by reconfiguring their IRS-user channels individually. Third, we assume perfect CSI on all links to focus on the performance analysis and optimization in this paper. However, the coexistence of the IRS and MAs may significantly increase the CSI estimation overhead. It is thus interesting to study whether these two technologies can be reciprocal in terms of channel estimation. Last but not least, other more general system setups, e.g., MIMO and/or multi-reflection-IRS systems, can also be studied to explore the interactions between MAs and multiple IRSs.